# Regret, Uncertainty, and Bounded Rationality in Norm-Driven Decisions

### v1

@@ Abstract
This study introduces an agent-based model that explains how regret, uncertainty, and social norms interact to shape vaccination behavior during epidemics. The model integrates three behavioral mechanisms—anticipated regret, evolving norms, and uncertainty-dependent trust—within a unified learning framework. Grounded in psychology and behavioral economics, it captures how individuals make probabilistic choices influenced by material payoffs, fear, trust, and social approval. Simulations of the Susceptible–Infected–Recovered process show that collective outcomes are best when agents display an intermediate level of rationality: they deliberate enough to respond to risk but remain flexible enough to adapt, avoiding the instability of both random and overly rigid decision-making. Regret exerts a dual influence—moderate levels encourage adaptive self-correction, while excessive regret or greed destabilize choices. Uncertainty has a similarly non-linear effect: moderate ambiguity promotes caution, but too much uncertainty disrupts coordination. Social norms restore cooperation by compensating for incomplete information. Among them, personal norms drive behavior when individuals have clear information and moral confidence; injunctive norms—reflecting perceived approval—gain influence under uncertainty; and descriptive norms, based on observing others’ actions, serve as informational cues that guide behavior when direct knowledge is limited. Overall, the model provides a psychologically grounded, computationally explicit account of how emotion, cognition, and social norms jointly govern preventive behavior during epidemics.
@@ 1 Introduction
Social norms—people’s perceptions of what others do (descriptive norms) and what others approve of (injunctive norms)—strongly shape vaccination behavior [3, 1, 4]. Descriptive norms guide behavior through informational cues and are most influential under uncertainty about others’ actions, whereas injunctive norms operate via moral and reputational channels [5]. Aligning these two norms enhances compliance, while misalignment undermines it [2, 6]. Large-scale studies show that providing accurate normative information—such as evidence of high vaccination rates—increases willingness to vaccinate, particularly among uncertain or mistrustful individuals [7]. What is more, evidence from experimental “pandemic-like” environments showed that individuals respond very differently to descriptive versus injunctive cues; in particular, Woike et al. [8] found that only injunctive messages consistently reduced risk-taking, while descriptive messages could potentially even lead to backfiring, i.e. increasing risk-taking behavior. These findings underscore the importance of treating social norms as multidimensional influences rather than uniform informational signals, highlight the need for models that capture how descriptive and injunctive expectations coevolve.
@@
Existing behavioral-epidemic models typically treat regret, uncertainty, and norms in isolation, limiting their explanatory power. This work unifies them in a single, psychologically grounded computational framework. It contributes (i) a dynamic weighting scheme linking emotional and informational uncertainty to social influence, (ii) an additive regret term capturing affective self-correction, and (iii) coupled evolution equations for three interdependent norm types. Together, these elements clarify how trust, emotion, and cognition coevolve under epidemic risk.
We address three central questions: (1) How do regret and uncertainty jointly shape vaccination decisions and epidemic outcomes? (2) Under what conditions do evolving social norms stabilize cooperation? (3) How does bounded rationality mediate the balance between self-correction and instability? By answering these, the model identifies when emotional and normative mechanisms promote resilience—or drive breakdown—in collective preventive behavior.
@@ 2 Model
Refer to captionRefer to caption
Fig 1: Left: Algorithm of system dynamics. The flowchart outlines the sequence of processes within each epidemic season. The model first simulates infection dynamics on the physical layer, estimating each agent’s infection probability. After each season, agents decide whether to vaccinate based on the previous outcomes, integrating both material (learning-based) and normative considerations. Subsequently, they update their norms before the next season begins. Top Right: Schematic of the model’s algorithm. The schematic summarizes the iterative process linking epidemic spread, vaccination decisions, and norm updates across seasons. Bottom Right: Dynamics of the vaccination game on a two-layer network. The model operates on two interconnected layers. On the physical layer, infection propagates through a Susceptible–Infected–Recovered (SIR) process; on the social layer, agents observe outcomes and peers’ actions to guide vaccination. The physical layer follows a small-world topology, while the social layer uses a Klimek–Thurner structure—both empirically supported as realistic for contact and communication networks. A significant overlap between layers reflects that many physical ties also transmit social information.
@@ 2.1.2 Uncertainty
Refer to caption Refer to caption
Fig 2: Left: Schematic of the decision-making process. The diagram illustrates how learning and social norms jointly determine each agent’s vaccination intention, weighted by their relative influence. Two assumptions guide this interaction: (a) The less safe an agent feels—i.e., the lower $S_{i}(t)$, representing infection risk from the previous season—the more they rely on empirical cues (learning or empirical expectations) rather than normative ones (personal beliefs or injunctive expectations). (b) The more heterogeneous and unstable the environment—when agents’ recent decisions diverge from one another or from past averages—the greater the reliance on personal experience and beliefs instead of social imitation. Right: Schematic of the dynamics of social norm variables. The diagram shows how different factors shape norm evolution under two key assumptions: (a) Each agent updates her norms with probability $\gamma_{i}$ under external influence, otherwise combining personal beliefs, existing norms, and peers’ actions. (b) Environmental instability—large deviations between recent and past behaviors—increases dependence on personal factors over conformity to others.
@@ 2.2 Experience dependent mechanism
2.2.1 Learning from Experience
@@
Although normative regret (e.g., guilt or social disapproval) could be modeled similarly, evidence indicates these emotions operate mainly ex post, reinforcing self-correction and social alignment rather than foresight [50, 53]. Accordingly, we restrict regret here to material outcomes and introduce normative regret later in the section on norm dynamics.
@@ 3.1 Social norms evolution
Figures 3A and 3B show the joint evolution of epidemic and behavioral dynamics under varying cognitive and normative conditions. In Fig. 3, infection levels rise with infectivity $\beta$, but both longer memory and social norms markedly reduce outbreak size. Agents with extended memory ($m=4$ vs. 1–2) incorporate more payoff history, vaccinate more consistently, and face smaller epidemics. Adding social norms (red line) further suppresses infections for the same memory length, even when average vaccination coverage changes little. Norms therefore improve the distribution rather than the level of vaccination—reducing clustering and stabilizing uptake. This pattern matches previous results [18], where endogenous norms sustained moderate coverage at low $\beta$ by preserving attitudinal diversity.
@@ 4 Conclusions
Methodologically, the Regret–Uncertainty model illustrates how agent-based modeling (ABM) can bridge psychological realism and epidemiological dynamics. By embedding cognitive, emotional, and normative heterogeneity into decision rules, it extends previous vaccination ABMs [66, 70, 22, 18] and generalizes experience-weighted attraction learning to include affective feedback and endogenous norm evolution. Unlike imitation models that treat norms as external, this framework internalizes moral and descriptive expectations, enabling targeted interventions and linking behavioral economics with computational epidemiology.

### v2

@@ Abstract
This study introduces an agent-based model to study how regret, uncertainty, and social norms interact to shape vaccination behavior during epidemics. The model integrates three behavioral mechanisms—anticipated regret, evolving norms, and uncertainty-de-pendent trust—within a unified learning framework. Grounded in psychology and behavioral economics, it captures how individuals make probabilistic choices influenced by material payoffs, fear, trust, and social approval. Simulations of the Susceptible–Infected–Recovered process show that collective outcomes are best when agents display an intermediate level of rationality: they deliberate enough to respond to risk but remain flexible enough to adapt, avoiding the instability of both random and overly rigid decision-making. Regret exerts a dual influence—moderate levels encourage adaptive self-correction, while excessive regret or greed destabilize choices. Uncertainty has a similarly non-linear effect: moderate ambiguity promotes caution, but too much uncertainty disrupts coordination. Social norms restore cooperation by compensating for incomplete information. Personal norms guide behavior when individuals have reliable information and feel confident in their judgments. Injunctive norms—signals of others’ approval—become more influential under uncertainty, while descriptive norms, which arise from observing others’ actions, provide informational cues that help people decide what to do when direct knowledge is limited. Overall, the model provides a psychologically grounded, computationally explicit account of how emotion, cognition, and social norms jointly govern preventive behavior during epidemics.
@@ 1 Introduction
Social norms—people’s perceptions of what others do (descriptive norms) and what others approve of (injunctive norms)—strongly shape vaccination behavior [3, 1, 4]. Descriptive norms guide behavior through informational cues and are most influential under uncertainty about others’ actions, whereas injunctive norms operate via moral and reputational channels [5]. Aligning these two norms enhances compliance, while misalignment undermines it [2, 6]. Large-scale studies show that providing accurate normative information—such as evidence of high vaccination rates—increases willingness to vaccinate, particularly among uncertain or mistrustful individuals [7]. What is more, evidence from experimental “pandemic-like” environments showed that individuals respond very differently to descriptive versus injunctive cues; in particular, Woike et al. [8] found that only injunctive messages consistently reduced risk-taking, while descriptive messages could potentially even lead to backfiring, i.e. increasing risk-taking behavior. These findings underscore the importance of treating social norms as multidimensional influences rather than uniform informational signals, highlighting the need for models that capture how descriptive and injunctive expectations coevolve.
@@
Existing behavioral-epidemic models typically treat regret, uncertainty, and norms in isolation, limiting their explanatory power. This work unifies them in a single, psychologically grounded computational framework. It contributes (i) a dynamic weighting scheme linking emotional and informational uncertainty to social influence, (ii) an additive regret term capturing affective self-correction, and (iii) coupled evolution equations for three interdependent norm types. Together, these elements contribute in clarifying how trust, emotion, and cognition coevolve under epidemic risk.
We address three central questions: (1) How do regret and uncertainty jointly shape vaccination decisions and epidemic outcomes? (2) Under what conditions do evolving social norms improve cooperation? (3) How does bounded rationality mediate the balance between self-correction and instability? By answering these, the model identifies when emotional and normative mechanisms promote resilience—or drive breakdown—in collective preventive behavior.
@@ 2 Model
Fig 1: Left: Algorithm of system’s dynamics. The flowchart outlines the sequence of processes within each epidemic season. The model first simulates infection dynamics on the physical layer, estimating each agent’s infection probability. After each season, agents decide whether to vaccinate based on the previous outcomes, integrating both material (learning-based) and normative considerations. Subsequently, they update their norms before the next season begins. Right (A): Schematic of the two layer multiplex network. The model operates on two interconnected layers. On the physical layer, infection propagates through a Susceptible–Infected–Recovered (SIR) process; on the social layer, agents observe outcomes and peers’ actions to guide vaccination. The physical layer follows a small-world topology, while the social layer uses a Klimek–Thurner structure—both empirically supported as realistic for contact and communication networks. A significant overlap between layers reflects that many physical ties also transmit social information. (B): Dynamics of the vaccination game. Schematic of the algorithm. Each season, the SIR dynamics run on the physical layer to estimate infection probabilities. Agents then decide whether to vaccinate, considering both payoff-based learning and normative factors, update their norms accordingly, and begin the next season.
@@ 2.1.2 Uncertainty
Fig 2: (A): Schematic of the decision-making process. The diagram illustrates how learning and social norms jointly determine each agent’s vaccination intention, weighted by their relative influence. Two assumptions guide this interaction: (a) The less safe an agent feels—i.e., the lower $S_{i}(t)$, representing infection risk from the previous season—the more they rely on empirical cues (learning or empirical expectations) rather than normative ones (personal beliefs or injunctive expectations). Similarly, the more uncertain the agent feels about the knowledge he posses, the more he will consult his injunctive, moral variables (personal and normative expectations). (b) The more heterogeneous and unstable the environment—when agents’ recent decisions diverge from one another or from past averages—the greater the reliance on personal experience and beliefs instead of social imitation. (B): Schematic of the dynamics of social norm variables. The diagram shows how different factors shape norm evolution under two key assumptions: (a) Each agent updates her norms with probability $\gamma_{i}$ under external influence, otherwise combining personal beliefs, existing norms, and peers’ actions. (b) Environmental instability—large deviations between recent and past behaviors and low consensus among neighbors—increases dependence on personal factors over conformity to others.
@@ 2.2 Experience dependent mechanism
Human behaviour reflects a mix of model-free and model-based reinforcement learning [24], combining accumulated experience with forward-looking considerations. The Experience-Weighted Attraction (EWA) framework captures this integration [46], bringing together reinforcement learning [47] and belief learning [48] by assigning equal weight to realized and forgone payoffs. In our setting, forgone payoffs are inferred from neighbours’ average infection rates because individuals do not observe others’ outcomes directly [49]. Agents evaluate payoffs over the past $m$ cycles, with memory decaying at a rate determined by perceived epidemiological risk. EWA thus describes how individuals merge past experience with beliefs about others’ behaviour. Its empirical support [50] makes it a suitable, cognitively grounded foundation for modelling social adaptation and norm compliance.
@@
Although normative regret (e.g., guilt or social disapproval) could be modeled similarly, evidence indicates these emotions operate mainly ex post, reinforcing self-correction and social alignment rather than foresight [55, 58]. Accordingly, we restrict regret here to material outcomes and introduce normative regret, i.e. social disapproval, in the next section on norm dynamics.
@@ 3.1 Social norms evolution
Figures 3A and 3B show the joint evolution of epidemic and behavioral dynamics under varying cognitive and normative conditions. In Fig. 3, infection levels rise with infectivity $\beta$, but both longer memory and social norms markedly reduce outbreak size. Agents with extended memory ($m=4$ vs. 1–2) incorporate more payoff history, vaccinate more consistently, and face smaller epidemics. Adding social norms (red line) further suppresses infections for the same memory length, even when average vaccination coverage changes little. Norms therefore improve the distribution rather than the level of vaccination—reducing clustering and stabilizing uptake.
@@ 4 Conclusions
Methodologically, the Regret–Uncertainty model illustrates how agent-based modeling can bridge psychological realism and epidemiological dynamics. By embedding cognitive, emotional, and normative heterogeneity into decision rules, it extends previous vaccination ABMs [71, 75, 22, 18] and generalizes experience-weighted attraction learning to include affective feedback and endogenous norm evolution. Unlike imitation models that treat norms as external, this framework internalizes moral and descriptive expectations, enabling targeted interventions and linking behavioral economics with computational epidemiology.